A steel wire of diameter 1 mm can support a tension of 0.2 kN. A steel cable to support a tension of 20 kN should have a diameter of what order of magnitude?
1
Expert's answer
2020-10-27T08:23:08-0400

The young modulus formula is given by

"\gamma=\frac{\textsf{Longitudinal stress}}{\textsf{Longitudinal strain}}\hspace{1cm}(1)"

Where "\gamma" is the young modulus

"\begin{aligned}\textsf{Longitudinal stress}=\frac{F}{A}\hspace{1cm}&(2)\\\textsf{Longitudinal strain}=\frac{∆L}{L}\hspace{0.9cm}&(3)\end{aligned}" Where "F,A,∆L\textsf{ and }L" represent the Force, Area, change in the length and the initial length respectively.

Substituting (2) and (3) in (1),

"\gamma=\frac{\frac{F}{A}}{\frac{∆L}{L}}\\
\gamma=\frac{FL}{∆LA}\\
A=\frac{FL}{\gamma∆L}"

Let "A_1" and "A_2" represent the cross sectional areas of the steel wire and steel cable respectively.

Let "F_1" and "F_2" represent the forces on the steel wire and steel cable respectively.

"A_1=\frac{F_1L}{\gamma\Delta L}\hspace{1cm}(4)\\
A_2=\frac{F_2L}{\gamma\Delta L}\hspace{1cm}(5)"

Dividing (5) by (4)

"\frac{A_2}{A_1}=\frac{\frac{F_2L}{\gamma\Delta L}}{\frac{F_1L}{\gamma\Delta L}}\\
\frac{A_2}{A_1}=\frac{F_2}{F_1}\hspace{1.3cm}(6)\\"

"A_1=\frac{\Pi {D_1}²}{4}\hspace{1cm}(7)\\
A_2=\frac{\Pi{D_2}²}{4}\hspace{1cm}(8)"

Where "D_1" and "D_2" represent the diameters of the steel cable and steel wire respectively.

Substituting (7) and (8) in (6),

"\frac{\frac{\Pi{D_2}²}{4}}{\frac{\Pi {D_1}²}{4}}=\frac{F_2}{F_1}"

"\frac{{D_2}²}{{D_1}²}=\frac{F_2}{F_1}\\
{D_2}²={D_1}²×\frac{F_2}{F_1}\\
{D_2}²=(1×10^{-3}m)²×(\frac{20kN}{0.2kN})\\
{D_2}²=(10^{-6}m²)×(100)\\
{D_2}²=10^{-4}m²\\
D_2=\sqrt{10^{-4}m²}\\
D_2=10^{-2}m\\
D_2=10mm"